OpenStudy (anonymous):

What is the product of the rational expressions below? x + 5/x + 7 * 10x / x - 7

OpenStudy (ghostgate):

So you have: (x + 5)/(x + 7) * (10x)/(x - 7) correct?

OpenStudy (anonymous):

yes

OpenStudy (ghostgate):

First things first, I would probably try to even out your denominator like you would with any ordinary fraction problem so, multiply the second fraction by '-1' first to even your denominators. (x + 5)/(x + 7) * -1(10x)/-1(x - 7) ~ Which gives you: (x + 5)/(x + 7) * -10x/-x + 7 ~ Hmmm... I thought that would work... hold on.

OpenStudy (anonymous):

10x • (x + 5) —————— but need this simplified (x + 7) • (x - 7)

OpenStudy (ghostgate):

Oh... thanks that actually helps a lot. Okay: 10x(x + 5) = 10x^2 + 50x --------- = (x + 7)(x - 7) = x^2 - 49 Your final answer being: 10x^2 + 50x ----------- x^2 - 49 I don't think you can simplify further, you do know how I got the (x+7)(x-7) to come out with x^2 - 49, right?